How much energy is lost by a 30.0g sample of iron that decreases in temperature from 56.7C to 25.0C? What if instead of iron it was a sample of water ?

SHOW WORK!!!!!
Answer #1

1) For iron

Given:

m = 30 g

C = 0.45 J/g.oC

Ti = 56.7 oC

Tf = 25 oC

use:

Q = m*C*(Tf-Ti)

Q = 30.0*0.45*(25.0-56.7)

Q = -428 J

Answer: 428 J

2) for water:

Given:

m = 30 g

C = 4.184 J/g.oC

Ti = 56.7 oC

Tf = 25 oC

use:

Q = m*C*(Tf-Ti)

Q = 30.0*4.184*(25.0-56.7)

Q = -3979 J

Answer: -3.98*10^3 J